# Study of Flat Spectrum Radio Quasars and BL Lacertae Objects as Sources of Diffusive Ultra High-Energy Cosmic Rays

### v1

@@ Abstract
We examine whether Flat Spectrum Radio Quasars (FSRQs) and BL Lacertae objects (BL Lacs) can act as plausible astrophysical sources of diffuse ultra-high-energy cosmic rays (UHECRs). Using realistic luminosity-dependent density evolution (LDDE) functions derived from observed gamma-ray luminosity functions for FSRQs and BL Lacs, we calculate the redshift evolution of the cosmic ray source population through integrated luminosity functions. The diffuse UHECRs flux from these sources is modelled by propagating nuclei through extragalactic space, including energy losses from interactions with cosmic photon backgrounds. The resulting UHECRs spectra are compared with observational data from the Pierre Auger Observatory and the Telescope Array, with fluxes normalised at reference energies. In addition to spectral comparisons, we generate HEALPix-based sky maps and identify potential hotspots. Near these hotspots, several galaxies and extragalactic sources are found, suggesting possible associations with the enhanced flux regions. Our results indicate that LDDE-modelled AGNs could contribute significantly to the observed diffuse UHECRs flux and provide constraints on their role as dominant sources.
@@ II.1 Flat Spectrum Radio Quasars (FSRQs)
In the PLE model, the comoving luminosity function $\Phi(L_{\gamma},z)$ is assumed to evolve only in luminosity, such that the redshift dependence enters through the evolution factor $e(z)$. Here, $L_{\gamma}$ denotes the $\gamma$-ray luminosity of the source and $\Phi(L_{\gamma},z)$ gives the number density of sources per unit luminosity. The local luminosity function at $z=0$ is characterized by a broken power-law with a normalization factor $A$, break luminosity $L_{*}$, low-luminosity slope $\gamma_{1}$, and high-luminosity slope $\gamma_{2}$, where $\gamma_{1}$ controls the behavior of the faint-end population ($L_{\gamma}\ll L_{*}$) and $\gamma_{2}$ determines the fall-off at the bright end ($L_{\gamma}\gg L_{*}$). The luminosity evolution factor, $e(z)=(1+z)^{k}\exp(z/\xi)$, describes how the characteristic luminosity shifts with redshift, where $k$ sets the strength of luminosity evolution and $\xi$ controls the redshift scale at which the evolution transits or saturates. As a simple PLE function model does not adequately reproduce the Fermi-LAT data, and given evidence for a luminosity-dependent evolution of the redshift peak, we fit the Fermi-LAT FSRQ sample using a luminosity-dependent density evolution (LDDE) model. In this approach, the evolution occurs mainly in density, with the redshift peak depending on luminosity. The LDDE model is parametrised as in Ref. [50], which is
@@ II.2 BL Lacertae (BL Lac) Objects
$\Phi(L_{\gamma},z=0,\Gamma)=\frac{A}{2.303\,L_{\gamma}}$ $\left[\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma_{1}}+\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma 2}\right]^{-1}\!\!\!\cdot$
$\exp\left[-\,\frac{(\Gamma-\mu(L_{\gamma}))^{2}}{2\sigma^{2}}\right]$ (9)
Here also, we utilise the same formalism as performed in Eqs. (4)–(6) and the best fit values of the parameters are adopted from Ref. [51].
@@ III Diffusion of Cosmic Rays in Turbulent Magnetic Fields
where $\gamma$ is the spectral index, and $a_{\text{I}}$ and $a_{\text{L}}$ are constants. For a Kolmogorov-type turbulence spectrum, $\gamma=5/3$ with $a_{\text{L}}\approx 0.23$ and $a_{\text{I}}\approx 0.9$. As noted earlier, the diffusion length $l_{\text{D}}$ represents the typical distance over which a particle’s total deflection becomes roughly one radian, defined as $l_{\text{D}}=3D/c$.
@@
Here, $H(t)E$ corresponds to adiabatic energy losses due to the expansion of the Universe, while $b_{\text{int}}(E)$ represents energy losses arising from interactions. UHECRs may interact with both the CMB and extragalactic background light (EBL). In this study, however, we consider only the energy losses caused by interactions with the CMB, which include processes such as pair production and photopion production [61].
@@ IV Flux of UHECRs
where $z_{\text{max}}$ is the maximum redshift at which the source starts emitting CRs, and $E_{\text{g}}$ is the generation energy at redshift $z$ corresponding to energy $E$ at $z=0$.
The total source emissivity $\mathcal{N}$ is obtained by summing the charge-specific emissivities $\mathcal{N}_{\text{Z}}$ for different nuclei. The charge-specific emissivity follows a power-law with a rigidity cutoff $ZE_{\text{max}}$ and is given by [81]
@@
where $\varepsilon_{\text{Z}}$ denotes the relative contribution of nuclei with charge $Z$ to the CR flux, and $f(z)$ describes the evolution of source emissivity with redshift $z$. In our study, we improve the description of $f(z)$ by adopting the luminosity-dependent density evolution (LDDE) model for FSRQs derived by Ajello et al. [50]. Instead of using a simple phenomenological form, often taken as a power-law or broken power-law in $(1+z)$, we employ a physically motivated redshift evolution based on the observed FSRQ luminosity function. In this framework, the comoving source density is expressed as
@@
Equation (19) can be generalised to nuclei in terms of their rigidity. During photo-disintegration, the rigidity and Lorentz factor of the primary fragment are largely conserved, minimally affecting the diffusion of the particle. However, these processes introduce complications because the source term $\mathcal{N}$ refers to the primary nucleus responsible for the observed fragment, and establishing this connection is challenging due to the stochastic nature of disintegration. This issue was discussed in Ref. [81], and we extend this discussion within modified and alternative gravity frameworks [82, 83].
For multiple sources, we apply the propagation theorem [84] to sum contributions from all sources, which satisfies
@@
Figure 1: Fluxes of UHECRs and $X_{\text{max}}$ values for a mixed composition scenario from FSRQs (top panel) and BL Lac objects (bottom panel), obtained by using different source separation distances (see text) and varying relative abundances of nuclear species. The fluxes and $X_{\text{max}}$ values are compared with Auger and TA measurements.
For the mixed composition case, we draw Fig. 1 for the fluxes along with the mean depths of shower maximum $\langle X_{\text{max}}\rangle$ as predicted by our theoretical calculations. The nuclear abundance of p, He, N, Si and Fe is mentioned at the top of the plots. Since we use two different AGN source population, so we get two different separation distance in fitting the numerical results with Auger and TA data. The corresponding values of the source separation distances are $160$ kpc and $0.2$ Mpc for FSRQs and BL Lac objects, respectively. The depth of the shower maximum, $X_{\text{max}}$, is determined using a parametrization based on air shower physics [86, 87]. For a nucleus with mass number $A$ and energy $E$, it is expressed as [86, 87].
$X_{\text{max}}(E,A)=X_{0}+\nu\ln\left(\frac{E}{A}\right),$ (27)
@@
Refer to captionRefer to caption
Figure 2: UHECR flux skymaps from FSRQs (top panel) and BL Lac objects (bottom panel) for $10$ EeV and $50$ EeV energies.
@@
where $\text{J}^{i}_{\text{th}}$ are the theoretical values of flux obtained from our calculations and $\text{J}^{i}_{\text{obs}}$ are the observed flux values, which are obtained from the Auger and the TA experiment. $\sigma$ denotes the total uncertainty in each energy bin, incorporating both statistical and systematic errors. The corresponding reduced $\chi^{2}$ values with the combined Auger and TA data are $2.02$ and $2.43$ for FSRQs and BL Lac objects respectively.
Fig. 1 also displays the fluxes for a mixed composition of nuclei along with the corresponding $\langle X_{\text{max}}\rangles$ and $\sigma\left(X_{\text{max}}\right)$s at the redshift $z_{\text{max}}=1$. The parametrisations show deviations from a purely proton composition, and the influence of the astrophysical source model is also apparent. The observational data for $\langle X_{\text{max}}\rangles$ from Auger and TA are taken from Refs. [89, 90]. These datasets provide $\langle X_{\text{max}}\rangle$ values along with statistical errors and uncertainties. In the $\langle X_{\text{max}}\rangle$ and $\sigma\left(X_{\text{max}}\right)$ plots, the statistical uncertainties are shown as shaded bands. The good agreement between our results and the observational data supports the validity of the source models used in reproducing the CR spectra. For readability, we include Table 1, which summarizes the key parameters used in fitting the UHECR energy spectrum with the Auger and TA observational data for the FSRQ and BL Lac source classes.
Table 1: Comparison of parameters for FSRQs and BL Lac objects.
@@
To model the CRs flux distributions, we use the HEALPix [91] framework with a resolution of $N_{\text{side}}=256$, corresponding to $N_{\text{pix}}=12\times N_{\text{side}}^{2}$ pixels covering the entire sky. Source positions are distributed uniformly in right ascension (RA) and declination (Dec), where RA ranges from $0^{\circ}$ to $360^{\circ}$ and Dec from $90^{\circ}$ to $90^{\circ}$. These coordinates are converted into HEALPix angular coordinates $(\theta,\phi)$ using $\theta=90^{\circ}-\text{Dec}$ and $\phi=\text{RA}$. The HEALPix pixel indices are obtained with healpy.ang2pix. To construct the flux maps, predefined flux values are assigned to each pixel, summing contributions from all sources in the pixel. A Gaussian smoothing filter with an angular scale of $2^{\circ}$ is applied using healpy.smoothing to ensure spatial smoothness. The maps are normalised with a logarithmic stretch to enhance contrast, replacing zero-valued pixels with a small nonzero value to avoid numerical artefacts. For visualisation, Mollweide projections [91] with a logarithmic colour scale are used, where longitude and latitude markers assist interpretation. This approach provides an accurate and clear depiction of the modelled CRs flux distributions across the sky.
The skymaps in Fig. 2 illustrate the CRs flux distributions for uniform source distribution models of the considered astrophysical sources across HEALPix pixels, with random sampling of (RA, Dec) coordinates converted into HEALPix $(\theta,\phi)$ angles. Fig. 2 presents the full-sky distribution of UHECRs flux in Mollweide projection for two distinct classes of AGN sources, namely FSRQs and BL Lacertae (BL Lac) objects, at energies of $10$ EeV and $50$ EeV. The top panels correspond to the FSRQ case, while the bottom panels illustrate the BL Lac scenario. In each case, the flux intensity is represented on a logarithmic color scale, where the red regions denote comparatively lower flux values and the blue regions indicate higher flux intensities. The maps reveal distinct anisotropy patterns in the UHECRs flux distribution, particularly evident in the FSRQ case. At lower energies ($10$ EeV), the flux appears more diffuse, indicating a larger propagation horizon and contributions from a greater number of distant sources. However, at higher energies (50 EeV), the anisotropy becomes more pronounced, with localized enhancements of flux (blue regions) corresponding to the dominance of nearby, high-luminosity FSRQs. This behaviour arises from the increased energy losses during propagation, which restrict the contribution of distant sources and amplifies the directional dependence of the flux. In contrast, the BL Lac sources display a relatively smoother flux distribution at both energies, suggesting a weaker degree of anisotropy. This can be attributed to their lower intrinsic luminosity and more uniform spatial distribution, which leads to a more isotropic flux pattern. Overall, the comparison between FSRQ and BL Lac source classes highlights the dependence of UHECRs anisotropy on both the source type and the cosmic ray energy, emphasizing the combined effects of source luminosity function, spatial evolution, and propagation constraints in shaping the observed flux distribution.
@@
Table 2: Nearby galaxies and extragalactic sources within $1^{\circ}$ of the hotspot located at (RA = 163.42, Dec = $88.26^{\circ}$) for FSRQ case.
@@
Table 3: Nearby galaxies and extragalactic sources within $1^{\circ}$ of hotspot located at (RA = 235.59°, Dec = -88.45°) for BL Lac case.
@@
TYC 9530-1462-1 RG* $4.14\times 10^{-5}$ 0.925 259.49 87.90-87.90
To identify galaxies or AGNs in the vicinity of the detected hotspots, we performed a positional cross-matching using the astroquery.simbad module. The coordinates of each hotspot (RA and Dec) were used as input to query the SIMBAD astronomical database within a circular search region of radius $1^{\circ}$. A custom SIMBAD query was configured to retrieve essential information such as the main identifier (main_id), object type (otypes), and equatorial coordinates (RA, Dec). The returned catalogue entries were then filtered to include only extragalactic sources, specifically those classified as galaxies or active nuclei. The following SIMBAD object types were considered: G (Galaxy), AGN, Radio (Radio Source), Sy (Seyfert Galaxy), and LINER (Low-Ionization Nuclear Emission-line Region). For each candidate object, the angular separation from the hotspot position was calculated using the astropy.coordinates.SkyCoord module, and only those lying within $1^{\circ}$ were considered. This method ensures the identification of nearby galaxies or AGN-like sources that could be potential counterparts or contributors to the observed hotspot regions in the sky. Numerous entries are classified as RG* (Red Giant Stars), and some are AG? (possible active galaxy) and GrG (galaxy group).
For FSRQs, using the above method, we searched for extragalactic sources within an angular radius of $1^{\circ}$ around the hotspot located at $(\mathrm{RA}=163.42^{\circ},\,\mathrm{Dec}=-88.26^{\circ})$. Notable entries from ESO 1-5, CPD, TYC, and UCAC are found in this region. The presence of nearby galaxies in the vicinity of this hotspot may suggest possible associations between these extragalactic sources and the observed high-energy CRs flux enhancement in this direction. For the BL Lac objects, among the identified sources within $1^{\circ}$ of the hotspot located at RA $235.59^{\circ}$, Dec $88.45^{\circ}$, several prominent extragalactic objects were found. A few possible active galaxies (AG?) like ESO 1-6 and ESO 1-7 were identified, suggesting the presence of AGNs in the vicinity. The catalogue also includes several RG* from CPD and TYC catalogue. Overall, the spatial correlation indicates that the region is populated with a mix of normal and active galaxies, as well as nearby stellar contaminants. For both FSRQs and BL Lac objects, these nearby galaxies or extragalactic sources are given in Tables 2 and 3.

### v2

@@ Abstract
We examine whether Flat Spectrum Radio Quasars (FSRQs) and BL Lacertae objects (BL Lacs) can act as plausible astrophysical sources of diffuse ultra-high-energy cosmic rays (UHECRs). Using realistic luminosity-dependent density evolution (LDDE) functions derived from observed gamma-ray luminosity functions for FSRQs and BL Lacs, we calculate the redshift evolution of the cosmic ray source population through integrated luminosity functions. The diffuse UHECRs flux from these sources is modelled by propagating nuclei through extragalactic space, including energy losses from interactions with cosmic photon backgrounds. The resulting UHECRs spectra are compared with observational data from the Pierre Auger Observatory and the Telescope Array, with fluxes normalised at reference energies. We also construct illustrative full skymaps of the integral UHECRs flux from $10$ EeV to $100$ EeV using the HEALPix framework, assuming a uniform distribution of synthetic sources and including energy-dependent magnetic diffusion. The regions of higher flux in the maps arise naturally from propagation effects and the limited UHECRs horizon. Our results indicate that LDDE-modelled AGNs could contribute significantly to the observed diffuse UHECRs flux and provide constraints on their role as potential sources.
@@ II.1 Flat Spectrum Radio Quasars (FSRQs)
In the PLE model, the comoving luminosity function $\Phi(L_{\gamma},z)$ is assumed to evolve only in luminosity, such that the redshift dependence enters through the evolution factor $e(z)$. Here, $L_{\gamma}$ denotes the $\gamma$-ray luminosity of the source and $\Phi(L_{\gamma},z)$ gives the number density of sources per unit luminosity. It is to be noted that the observed $\gamma$-ray luminosities of blazars are affected by Doppler boosting and hence $L_{\gamma}$ is related with the intrinsic luminosity $L_{\gamma}^{\prime}$ as $L_{\gamma}=(\delta_{\rm e}^{6}/\Gamma_{\rm e}^{2})L_{\gamma}^{\prime}$ [50, 51], where $\delta_{\rm e}$ is the Doppler factor of the jet, which quantifies the enhancement of the observed luminosity $L_{\gamma}$ due to relativistic motion, and $\Gamma_{\rm e}$ is the Lorentz factor. The local luminosity function at $z=0$ is characterized by a broken power-law with a normalization factor $A$, break luminosity $L_{*}$, low-luminosity slope $\gamma_{1}$, and high-luminosity slope $\gamma_{2}$, where $\gamma_{1}$ controls the behavior of the faint-end population ($L_{\gamma}\ll L_{*}$) and $\gamma_{2}$ determines the fall-off at the bright end ($L_{\gamma}\gg L_{*}$). The luminosity evolution factor, $e(z)=(1+z)^{k}\exp(z/\xi)$, describes how the characteristic luminosity shifts with redshift, where $k$ sets the strength of luminosity evolution and $\xi$ controls the redshift scale at which the evolution transits or saturates. As a simple PLE function model does not adequately reproduce the Fermi-LAT data, and given evidence for a luminosity-dependent evolution of the redshift peak, we fit the Fermi-LAT FSRQ sample using a luminosity-dependent density evolution (LDDE) model. In this approach, the evolution occurs mainly in density, with the redshift peak depending on luminosity. The LDDE model is parametrised as in Ref. [50], which is
@@ II.2 BL Lacertae (BL Lac) Objects
$\Phi(L_{\gamma},z=0,\Gamma)=\frac{A}{2.303\,L_{\gamma}}$ $\left[\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma_{1}}+\left(\frac{L_{\gamma}}{L_{*}}\right)^{\gamma 2}\right]^{-1}$
$\times\exp\left[-\,\frac{(\Gamma-\mu(L_{\gamma}))^{2}}{2\sigma^{2}}\right].$ (9)
As in the FSRQs case, in this case the relation between the observed luminosity $L_{\gamma}$ and the intrinsic luminosity $L_{\gamma}^{\prime}$ is $L_{\gamma}=\delta_{\rm e}^{4}L_{\gamma}^{\prime}$ [52, 51]. Also, we utilise here the same formalism as performed in Eqs. (4)–(6) and the best fit values of the parameters are adopted from Ref. [52].
@@ III Diffusion of Cosmic Rays in Turbulent Magnetic Fields
where $\gamma$ is the spectral index, and $a_{\text{I}}$ and $a_{\text{L}}$ are constants. For a Kolmogorov-type turbulence spectrum, $\gamma=5/3$ with $a_{\text{L}}\approx 0.23$ and $a_{\text{I}}\approx 0.9$. As noted earlier, the diffusion length $l_{\text{D}}$ represents the typical distance over which a particle’s total deflection becomes roughly one radian, defined as $l_{\text{D}}=3D/c$. Throughout this work, we assume a homogeneous turbulent intergalactic magnetic field with RMS strength $B_{\rm rms}=1~\mathrm{nG}$ and coherence length $l_{\rm c}=0.1~\mathrm{Mpc}$, corresponding to the maximum scale of a Kolmogorov turbulence spectrum.
@@
Here, $H(t)E$ corresponds to adiabatic energy losses due to the expansion of the Universe, while $b_{\text{int}}(E)$ represents energy losses arising from interactions. In this work, all interaction losses are calculated using only the CMB as the target photon field. For proton primaries, we include energy losses due to pair production and photopion production upon interactions with the CMB, treated in the continuous energy-loss approximation [62]. For nuclear primaries ($A>1$), we include pair production and photodisintegration caused by interactions with the CMB only; photon fields from the extragalactic background light (EBL) are not included. This approximation is appropriate for the class of sources considered here, which are able to accelerate nuclei to UHEs without strong photodisintegration [64, 65]. According to the Ref. [66], photodisintegration largely preserves the Lorentz factor and rigidity of the main nuclear fragment. As a result, this process does not change the diffusion coefficient or the overall form of particle propagation, and its effect enters only through the nuclear survival probability. Similar treatments have also been discussed in the context of modified and alternative gravity models [67, 68, 19].
@@ IV Flux of UHECRs
where $z_{\text{max}}$ is the maximum redshift at which the source starts emitting CRs, and $E_{\text{g}}$ is the generation energy at redshift $z$ corresponding to energy $E$ at $z=0$. The total source emissivity $\mathcal{N}$ is obtained by summing the charge-specific emissivities $\mathcal{N}_{\text{Z}}$ for different nuclei. The charge-specific emissivity follows a power-law with a rigidity cutoff $ZE_{\text{max}}$ and is given by [66]
@@
where $\varepsilon_{\text{Z}}$ denotes the relative contribution of nuclei with charge $Z$ to the CR flux, and $f(z)$ describes the evolution of source emissivity with redshift $z$. In our study, we improve the description of $f(z)$ by adopting the LDDE model for FSRQs and BL Lacs derived by Ajello et al. [50, 52]. Instead of using a simple phenomenological form, often taken as a power-law or broken power-law in $(1+z)$, we employ a physically motivated redshift evolution based on the observed FSRQ luminosity function. In this framework, the comoving source density is expressed as
@@
Equation (19) can be generalised to nuclei in terms of their rigidity [66]. For multiple sources, we apply the propagation theorem [87] to sum contributions from all sources, which satisfies
@@
Figure 1: Fluxes of UHECRs and $X_{\text{max}}$ values for a mixed composition scenario from FSRQs (top panel) and BL Lac objects (bottom panel), obtained by using different source separation distances (see text) and varying relative abundances of nuclear species. The observational values of flux of these two experiments are taken from [89] and those for $X_{\text{max}}$ are from [90, 91].
For the mixed composition case, we draw Fig. 1 for the fluxes along with the mean depths of shower maximum $\langle X_{\text{max}}\rangle$ as predicted by our theoretical calculations. The nuclear abundance of p, He, N, Si and Fe is mentioned at the top of the plots. Since we use two different AGN source population, so we get two different separation distance in fitting the numerical results with Auger and TA data [89]. The corresponding values of the source separation distances are $160$ kpc and $36$ Mpc for FSRQs and BL Lac objects, respectively. The depth of the shower maximum, $X_{\text{max}}$, is determined using a parametrization based on air shower physics [92, 93]. For a nucleus with mass number $A$ and energy $E$, it is expressed as [92, 93].
$X_{\text{max}}(E,A)=X_{0}+\nu~\text{log}_{10}\left(\frac{E}{A}\right),$ (27)
@@
Figure 2: Mollweide skymaps of the integral UHECRs flux in the range of $10\leq E\leq 100~\mathrm{EeV}$ for synthetic FSRQ (left) and BL Lac (right) populations. The maps assume a uniform comoving distribution of hypothetical sources, with no observational catalogues used. A $2^{\circ}$ Gaussian smoothing is applied. The colour scale shows $E^{3}J(E)$ in $\mathrm{eV^{2}\,m^{-2}\,s^{-1}\,sr^{-1}}$.
@@
where $\text{J}^{i}_{\text{th}}$ are the theoretical values of flux obtained from our calculations and $\text{J}^{i}_{\text{obs}}$ are the observed flux values, which are obtained from the Auger and the TA experiment. $\sigma$ denotes the total uncertainty in each energy bin, incorporating both statistical and systematic errors. The corresponding reduced $\chi^{2}$ values with the combined Auger and TA data are $2.41$ and $2.63$ for FSRQs and BL Lac objects respectively.
Fig. 1 also displays the fluxes for a mixed composition of nuclei along with the corresponding $\langle X_{\text{max}}\rangles$ and $\sigma\left(X_{\text{max}}\right)$s at the redshift $z_{\text{max}}=1$. The parametrisations show deviations from a purely proton composition, and the influence of the astrophysical source model is also apparent. The observational data for $\langle X_{\text{max}}\rangles$ from Auger and TA are taken from Refs. [90, 91]. These datasets provide $\langle X_{\text{max}}\rangle$ values along with statistical and systematic errors. The red and blue bands in $\langle X_{\text{max}}\rangles$ and $\sigma\left(X_{\text{max}}\right)$s plots denote the systematic errors of Auger and TA data, respectively. In the $\langle X_{\text{max}}\rangle$ and $\sigma\left(X_{\text{max}}\right)$ plots, the statistical uncertainties are shown as shaded bands. The good agreement between our results and the observational data supports the validity of the source models used in reproducing the CR spectra. For readability, we include Table 1, which summarizes the key parameters used in fitting the UHECR energy spectrum with the Auger and TA observational data for the FSRQ and BL Lac source classes.
It should be mentioned that the absolute $\langle X_{\max}\rangle$ values measured by the Auger and the TA should not be compared directly. As demonstrated in joint Auger-TA analyses, differences in detector design, event-selection criteria, and reconstruction methods introduce experiment-dependent offsets in $\langle X_{\max}\rangle$ that are unrelated to mass composition. Consequently, composition inference is based on the behaviour of $\langle X_{\max}\rangle$ and $\sigma(X_{\max})$ relative to single-species expectations and on the energy dependence of these observables, rather than on a direct comparison between these two datasets.
Table 1: Comparison of parameters for FSRQs and BL Lac objects. $\varepsilon$ denotes the relative contribution from the nuclei.
@@
$\varepsilon_{\text{p}}$ $0.5$ $0.45$
$\varepsilon_{\text{He}}$ $0.25$ $0.28$
$\varepsilon_{\text{N}}$ $0.15$ $0.15$
$\varepsilon_{\text{Si}}$ $0.06$ $0.07$
$\varepsilon_{\text{Fe}}$ $0.04$ $0.05$
To describe the angular distribution of the UHECRs flux, we generate full-sky intensity maps using the HEALPix framework [95] with a resolution parameter of ${\tt Nside}=256$, corresponding to $N_{\text{pix}}=12\times N_{\text{side}}^{2}$ pixels covering the full sky. In this work, all skymaps are built using a uniform comoving distribution of synthetic sources representing the FSRQ and BL Lac populations. No observational source catalogues are used. The positions of the simulated sources are generated randomly over the entire sky in right ascension ($0^{\circ}-360^{\circ}$) and declination ($90^{\circ}-+90^{\circ}$), and then converted to the corresponding HEALPix coordinates $(\theta,\phi)$. For each simulated source, we calculate its contribution to the integral UHECRs flux within the energy range $10\leq E\leq 100~\mathrm{EeV}$ shown in the maps. This is done by numerically integrating the propagated energy spectrum from the chosen energy threshold up to the maximum injection energy, and the resulting flux is added to the appropriate HEALPix pixel. Magnetic deflections are included through the diffusion coefficient $D(E)$, which depends on the strength and coherence length of the intergalactic magnetic field. These parameters in the Syrovatskii variable $\lambda^{2}(E,z)$ describe how the flux from each source is reduced with distance due to magnetic spreading during propagation. To reduce pixel-to-pixel numerical noise and to make large-scale anisotropy patterns easier to see, the final maps are smoothed using a Gaussian function with a width of $2^{\circ}$.
This construction is meant to be illustrative and is not intended to reproduce the observed UHECRs sky. Since the source population is uniformly distributed and entirely synthetic, any localized regions of higher flux (“hotspots”) arise naturally from statistical effects. In the energy range $10\leq E\leq 100~\mathrm{EeV}$, the propagation horizon is limited, so only a small number of nearby synthetic sources contribute significantly to the total flux. This produces apparent hotspots in the skymaps. These features, therefore, reflect the geometry of the simulated source distribution rather than real astrophysical anisotropies. Fig. 2 shows the resulting full-sky integral flux maps for the FSRQ and BL Lac scenarios. The FSRQ case exhibits stronger anisotropy, which is consistent with its higher typical luminosities and LDDE evolution, while the BL Lac case leads to a smoother flux distribution. The hotspots seen in the maps are therefore a natural outcome of the synthetic source model.
@@
Table 2: Nearby galaxies and extragalactic sources within $1^{\circ}$ of the hotspot located at (RA = 253.12, Dec = $87.08^{\circ}$) for FSRQ case.
@@
Table 3: Nearby galaxies and extragalactic sources within $1^{\circ}$ of hotspot located at (RA = 286.88, Dec = 87.08) for BL Lac case.
@@
To identify galaxies or AGNs in the vicinity of the detected hotspots, we performed a positional cross-matching using the astroquery.simbad module. The coordinates of each hotspot (RA and Dec) were used as input to query the SIMBAD astronomical database within a circular search region of radius $1^{\circ}$. Since the skymaps are produced entirely from synthetic sources that are uniformly distributed and therefore do not contain any observational information. Hence, any comparison with the positions of known AGNs from catalogues such as SIMBAD is made only for reference. This comparison is used simply to show which well-known extragalactic objects happen to lie close to the directions of enhanced flux in the model skymaps. However, the visual overlap between hotspots and catalogue objects indicates the efficiency of the predictability of the model considered here.
A custom SIMBAD query was configured to retrieve essential information such as the main identifier (main_id), object type (otypes), and equatorial coordinates (RA, Dec). The returned catalogue entries were then filtered to include only extragalactic sources, specifically those classified as galaxies or active nuclei. The following SIMBAD object types were considered: G (Galaxy), AGN, Radio (Radio Source), Sy (Seyfert Galaxy), and LINER (Low-Ionization Nuclear Emission-line Region). For each candidate object, the angular separation from the hotspot position was calculated using the astropy.coordinates.SkyCoord module, and only those lying within $1^{\circ}$ were considered. This method ensures the identification of nearby galaxies or AGN-like sources that could be potential counterparts or contributors to the observed hotspot regions in the sky. Numerous entries are classified as RG* (Red Giant Stars), and some are AG? (possible active galaxy) and GrG (galaxy group).
For FSRQs, using the above method, we searched for extragalactic sources within an angular radius of $1^{\circ}$ around the hotspot located at $(\mathrm{RA}=253.12^{\circ},\,\mathrm{Dec}=-87.08^{\circ})$. Notable entries from ESO 1-5, CPD, TYC, and UCAC are found in this region. The presence of nearby galaxies in the vicinity of this hotspot may suggest possible associations between these extragalactic sources and the observed high-energy CRs flux enhancement in this direction. For the BL Lac objects, among the identified sources within $1^{\circ}$ of the hotspot located at RA $286.88^{\circ}$, Dec $87.08^{\circ}$, several prominent extragalactic objects were found. A few possible active galaxies (AG?) like ESO 1-6 and ESO 1-7 were identified, suggesting the presence of AGNs in the vicinity. The catalogue also includes several RG* from CPD and TYC catalogue. Overall, the spatial correlation indicates that the region is populated with a mix of normal and active galaxies, as well as nearby stellar contaminants. For both FSRQs and BL Lac objects, these nearby galaxies or extragalactic sources are given in Tables 2 and 3.